OpenStudy (anonymous):

0.8n+1.22=0.408-0.6n

OpenStudy (anonymous):

0.8n + 0.6n = 0.408 - 1.22 1.4n = -0.812 n = -0.58

OpenStudy (anonymous):

What you want to do in this problem is isolate the variable "n". So just think of "n" like any other number. For example, take the statement 8+4=12 if the equation was 8 + n = 12 , we'd want to isolate n so what we would do is subtract 8 from both the left side and the right side and you would have n = 12-8 = 4. similarly, suppose you had a statement 2*3=6 if you were not given one of the numbers like in the previous example, and instead the equation was: 2n=6 you would divide 2 from both sides, thus isolating n and giving the equation n=3 Using these rules together, try to solve the equation by getting n all by itself.

OpenStudy (anonymous):

thanks alot 32p+7=9.56

OpenStudy (anonymous):

So in this equation, what do you want to have by the end?

OpenStudy (anonymous):

.8

OpenStudy (anonymous):

Yes, you got the hang of it. The answer is .08 but I assume it was a typo. What you would have done in this equation is try to isolate p by first subtracting 7 from both sides then you would get 32p=2.56 , then just divide both sides by 32 and you get p=.08

OpenStudy (anonymous):

3.5t+18.75=1.5-t

OpenStudy (anonymous):

2/5/90.8+1/3to 3rd power